OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

IM TOTES CLUELESS! #HELP

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Well think of 3 consecutive integers like 1,2,3 They are all one more than the last. So you can rewrite this: \((x)+(x+1)+(x+2)=267\)

OpenStudy (anonymous):

im not sure how 🙈

OpenStudy (midhun.madhu1987):

Let the 3 consecutive integers be x , x+1 and x + 2 Then their sum = 267 So, x + x + 1 + x + 2 = 267 3x + 3 = 267 Find x from this

OpenStudy (anonymous):

Solve for x in the equation

OpenStudy (anonymous):

do u times 3 times x???

jimthompson5910 (jim_thompson5910):

you have to solve 3x + 3 = 267

OpenStudy (anonymous):

i know but do it do 3 plus 3 then times x?

jimthompson5910 (jim_thompson5910):